1. 2 angles are supplementary. 1 angle is 5 more than 3 times the other. find the angles??

Found 2 solutions by jessica43, JBarnum:
Answer by jessica43(140)

To solve this problem, you are going to have to write two equations using what you know.
First, you know that you have two supplementary angles. This means that together the angles add up to 180 degrees. So:
A + B = 180 (where A = measure of angle one, B = measure of angle two)
Second, you now that angle one is 5 more than 3 times angle two:
A = 5 + 3B
Now you can plug in the second equation into the first and solve for B:
A + B = 180
(5 + 3B) + B = 180
5 + 4B = 180
4B = 175
B = 43.75
Finally, plug this value into the first equation to find A:
A + B = 180
A + 43.75 = 180
A = 136.25
So angle A is 136.25 degrees and angle B is 43.75 degrees.

Answer by JBarnum(2044)

supplementary means 2 angles that add up to 180 degrees
angle x and y
x+y=180
x=5+3y
5+3y+y=180
5+4y=180
4y=175
y=43.75
x=136.25